Three long, parallel conductors each carry a curre
Answer #1

a) at point A

Bnet = sqrt(2)*mue*I/(2*pi*a) + mue*I/(2*pi*3*a)

= sqrt(2)*4*pi*10^-7*1.94*1.94/(2*pi*0.007) + 4*pi*10^-7*1.94*1.94/(2*pi*3*0.007)

= 1.88*10^-4 T

= 188 micro T (downward)

b) Bnet = mue*I/(2*pi*2*a)

= 4*pi*10^-7*1.94*1.94/(2*pi*2*0.007)

= 5.38*10^-5 T

= 53.8 micro T (downward)

c) Bnet = sqrt(2)*mue*I/(2*pi*a) - mue*I/(2*pi*a)

= sqrt(2)*4*pi*10^-7*1.94*1.94/(2*pi*0.007) - 4*pi*10^-7*1.94*1.94/(2*pi*0.007)

= 4.45*10^-5 T

= 44.5 micro T (upward)